# HACD: Hand-Aware Conditional Diffusion for Monocular Hand-Held Object Reconstruction

### 23 Nov 2023 v1

@@ Abstract
Reconstructing hand-held objects from a single RGB image without known 3D object templates, category prior, or depth information is a vital yet challenging problem in computer vision. In contrast to prior works that utilize deterministic modeling paradigms, which make it hard to account for the uncertainties introduced by hand- and self-occlusion, we employ a probabilistic point cloud denoising diffusion model to tackle the above challenge. In this work, we present Hand-Aware Conditional Diffusion for monocular hand-held object reconstruction (HACD), modeling the hand-object interaction in two aspects. First, we introduce hand-aware conditioning to model hand-object interaction from both semantic and geometric perspectives. Specifically, a unified hand-object semantic embedding compensates for the 2D local feature deficiency induced by hand occlusion, and a hand articulation embedding further encodes the relationship between object vertices and hand joints. Second, we propose a hand-constrained centroid fixing scheme, which utilizes hand vertices priors to restrict the centroid deviation of partially denoised point cloud during diffusion and reverse process. Removing the centroid bias interference allows the diffusion models to focus on the reconstruction of shape, thus enhancing the stability and precision of local feature projection. Experiments on the synthetic ObMan dataset and two real-world datasets, HO3D and MOW, demonstrate our approach surpasses all existing methods by a large margin.
@@ 1 Introduction
Figure 1: We present hand-aware conditioning, utilizing both semantic and geometric embeddings to model hand-object interaction. We further propose hand-constrained centroid-fixed diffusion, where the centroid of the partially denoised point cloud at each step is constrained by the hand vertices and kept fixed throughout the whole diffusion and the reverse process.
@@
Modelling hand-object interactions has a wide range of applications across various domains, including AR/VR [38, 54, 55] and robotics [75, 24, 72]. Significant progress has been made in this task [38, 72, 4], in which single-view hand-held object reconstruction draws particularly great attention [30, 37, 73]. Reconstructing hand-held objects from a single RGB image is a highly challenging and ill-posed problem. Suffering from the lack of real-world data and the ambiguity caused by hand- and self-occlusion, the performance of single-view hand-held object reconstruction remains limited.
@@
Most existing works rely on Signed Distance Fields (SDFs) [37, 73] to represent the object shapes. Though effective under favorable conditions, such technique tends to result in over-smoothed and undetailed reconstructions [51, 10]. Moreover, prior works [30, 37, 73] typically utilize a deterministic modeling paradigm, which makes it hard to reason about the uncertainties introduced by hand- and self-occlusion. Recently, probabilistic point cloud denoising diffusion models [43, 80, 44] have shown to be effective for the task of single-view object reconstruction. First, compared to works employing surface-based representations like SDFs, approaches reconstructing point clouds enjoy better capabilities in overcoming artifacts such as noise, sparsity, and irregularity [10]. Second, the probabilistic nature of denoising diffusion models is particularly beneficial for modeling uncertainty and indeterminacy. Therefore, in this work, we propose to leverage these probabilistic point cloud denoising diffusion models to conduct hand-held object reconstruction from a single RGB image.
@@
Directly employing diffusion models in single-view hand-held object reconstruction, however, faces two main problems. First, diffusion models for 3D reconstruction typically assume isolated and unoccluded objects in the images [80, 44, 43] with self-occlusion properly modeled, while in hand-held object reconstruction tasks, hand occlusion needs to be modeled as well. In fact, the shape of the hand-held object is constrained by and highly related to the hand semantically and geometrically. Second, the object centroid is generally constrained by the hand vertices during the diffusion and the reverse processes. In existing diffusion models, the centroid of a partially denoised point cloud is not controlled and left variable, thus the centroid can easily deviate and move, for example, to the back side of the hand or even intersect with the hand, leading to physically implausible results. Additionally, the centroid bias causes the misalignment of the semantic features [19] and can thus have adverse effects on the object reconstruction quality.
@@
To solve the aforementioned problems, we present HACD, a Hand-Aware Conditional point cloud denoising Diffusion probabilistic model for single-view hand-held object reconstruction (Fig. 1). To fully utilize the hand prior, we introduce hand-aware conditioning to model hand-object interaction from both semantic and geometric perspectives. Via extraction of local features from object and hand, we acquire a unified hand-object semantic embedding, which compensates for the semantic feature deficiency induced by the hand-occluded object region. Further, at each step, we explicitly encode the relative position between hand articulation and the partially denoised point cloud to obtain a geometric hand articulation embedding. To fuse the aforementioned semantic and geometric embeddings, we propose a novel dual-stream conditioning to separately predict noise features in two branches and then estimate the final noise from the integrated feature. Additionally, a hand-constrained centroid fixing scheme is proposed to ensure that the centroid of the partially denoised point cloud does not diverge during the diffusion as well as the reverse process. In particular, we use a small and efficient neural network to estimate the hand-constrained object centroid to properly guide the reverse diffusion. As a consequence, the model only needs to take care of the simpler shape diffusion, rather than having to simultaneously account for shape and centroid. Moreover, leveraging the hand vertices constraint together with our centroid fixing can further enhance the precision and stability of our semantic and geometric features. Our contributions can be summarized as follows.
  • We present HACD, a hand-aware conditional point cloud diffusion model for monocular hand-held object reconstruction, demonstrating superior performance in modeling the uncertainties induced by hand- and self-occlusion.
  • We introduce a unified semantic hand-object embedding and a geometric hand articulation embedding to model semantic and geometric aspects of the shown hand-object interaction, which is subsequently utilized within our proposed dual-stream conditioning.
@@
  • We further propose a hand-constrained centroid fixing scheme, utilizing the hand vertices prior to prevent the centroid of the partially denoised point cloud from diverging during diffusion and its reverse processes.
@@
Experiments on the synthetic ObMan [30] dataset and two real-world datasets, namely HO3D [29] and MOW [4], demonstrate that HACD can surpass all existing methods by a large margin.
@@ 2 Related Work
Hand Pose estimation. Hand pose estimation methods from RGB(-D) images can be primarily categorized into model-free and model-based methods. Model-free approaches typically detect 2D keypoints, which are then lifted to 3D joint positions [57, 58, 82, 47, 50, 34, 48], while model-based approaches exploit statistical models like MANO [59] with low-dimensional parameter spaces [63, 1, 79, 81, 60]. Compared with model-free methods, the model-based methods exhibit better robustness to occlusions and domain discrepancies. Therefore, our method utilizes an off-the-shelf model-based approach [60] to predict hand poses, which are then further utilized for shape inference of hand-held objects.
Single-view Object Reconstruction. Object reconstruction is a long-standing problem in the computer vision community. Traditionally, varieties of works focus on reconstruction from multiple-view geometric techniques [61, 76, 15, 16]. Recently, advanced by the powerful representation capacity of deep learning, single-view-based methods for object reconstruction have shown promising results [44, 11, 70, 71], in spite of the highly ill-posed setting. Early learning-based methods propose to learn category-specific networks [36, 35, 67], while more recent works try to learn a generalizable model across multiple categories for either meshes [28, 26], voxels [25, 11, 69], point clouds [22, 39], or implicit representations such as NeRFs [45] and SDFs [51, 7]. In this work, we focus on the difficult yet very important problem of learning hand-held object reconstruction from a single RGB image [30], putting a particular emphasis on modeling the impact of the hand occlusion.
Hand-Held Object Reconstruction. Reconstructing hand-held objects is very challenging due to intricate mutual occlusions and self-occlusions. Earlier works opt for the simplified task of 6DoF object pose estimation [17, 64, 66, 29, 31], assuming known object templates. To better utilize hand-object interactions, some studies reason about hand and object poses by means of implicit feature fusion [6, 40, 62, 66], exploiting explicit geometric constraints [12, 2, 27, 4, 18, 77, 78], or enforcing of physical constraints [68, 52]. Increasing attention has recently been paid to model-free hand-held object reconstruction, as it is a more applicable setting. Exemplarily, while [30] explicitly predicts the object mesh, [37, 8, 9] leverage implicit 3D representations such as signed distance functions for reconstructing the shape of the hand and object. However, all these methods follow a deterministic approach, oftentimes resulting in low-quality reconstruction results for occluded and invisible parts, especially in the ill-posed single-view setting. We instead draw inspiration from recent advances in probabilistic 3D generation [33, 80, 19, 44] and additionally leverage hand-articulation priors to achieve higher-fidelity hand-held object reconstruction.
@@ 3 Method
Figure 2: Architecture of HACD. (I) Given a single-view RGB image, we first predict the hand pose $\phi_{H}$ and camera view $\phi_{C}$ by an off-the-shelf network. (II) The object centroid $\widehat{\mathcal{M}}$ is then estimated by our simple yet efficient hand-constrained centroid prediction network. (III) We further introduce a centroid-fixed diffusion network, which always keeps the centroid of partially denoised point cloud fixed at the predicted centroid $\widehat{\mathcal{M}}$ during the reverse process. (IV) Both the unified semantic hand-object embedding and geometric hand articulation embedding are utilized as condition in the diffusion and the reverse process, which is then fused by a dual-stream conditioning scheme.
Given a single-view RGB image $I$ capturing a hand-held object, we aim at reconstructing the 3D point cloud of the object. As shown in Fig. 2, we first utilize an off-the-shelf method [30, 60] to predict hand parameters $\phi_{H}$ and camera view $\phi_{C}$, where $\phi_{H}$ is defined using the MANO [59] hand model with 45DoF joint parameters and $\phi_{C}$ represents the 6DoF pose of the hand wrist in the world reference system. A hand-constrained centroid prediction network is then employed to predict the object centroid $\widehat{\mathcal{M}}$ given the estimated hand vertices. Eventually, we use our centroid-fixed conditional point cloud diffusion model to reconstruct the object point cloud. Compared with existing diffusion models, our model can thus fully concentrate on the reconstruction of the shape, leading to an improved performance. We further leverage semantic as well as geometric hand priors during diffusion. The unified semantic hand-object embedding and geometric hand articulation embedding are further used within our dual-stream conditioning. Due to the probabilistic nature of diffusion models together with our explicit modeling of hand-object interaction, HACD shows superior performance in handling uncertainties arising from hand- and self-occlusion.
In the following sections, we first introduce the fundamentals of conditional point cloud denoising diffusion models (Sec. 3.1). We then explain our novel hand-aware conditioning (Sec. 3.2), before describing our hand-constrained centroid-fixing scheme (Sec. 3.3).
@@ 3.1 Conditional Point Cloud Denoising Diffusion
We formulate single-view hand-held object reconstruction as conditional point cloud denoising diffusion, consisting of two Markov chains called the diffusion process and the reverse diffusion process. Suppose that we have a target point cloud with $N$ points $X_{0}\in\mathbb{R}^{3N}$ from the conditional distribution $q(X|z)$, where $z=I_{\phi_{C},\phi_{H}}$ is the input RGB image capturing the hand-held object with the corresponding camera view $\phi_{C}$ and the hand pose $\phi_{H}$. For the diffusion process, we gradually add Gaussian noise to the target point cloud at different levels $t\in\{1,\cdots,T\}$ as
@@
Starting with a point cloud sample $X_{T}$ from random Gaussian noise, the reverse process iteratively samples from $q(X_{t-1}|X_{t},z)$ to remove the added noise from the diffusion process. To approximate this reverse process, we train a conditional denoiser $\mathcal{D}_{\theta}(X_{t},t,z)$ to learn the distribution $q(X_{t-1}|X_{t},z)$ by
@@
3.2 Hand-Aware Conditioning
The essence of the conditional point cloud denoiser $\mathcal{D}_{\theta}(X_{t},t,z)$ is the modeling of the condition $z$ so to fully utilize the information of the input image along with the corresponding camera view and hand pose. Our key insight is that the hand can offer both semantic and geometric priors to facilitate object reconstruction. Therefore, a unified semantic hand-object embedding and a geometric hand articulation embedding are proposed in an effort to well model the hand-object interaction.
Unified Semantic Hand-Object Embedding. The 2D image feature is a vital component for object reconstruction. Unlike prior works that extract a global embedding for the image [11, 70, 71], we choose to compute unique semantic features for each point. Specifically, we first extract image features $\mathcal{F}\in\mathbb{R}^{H\times W\times C}$ by a standard 2D network such as ResNet [32] or ViT [21], with $C$ being the number of feature channels. Instead of sharing the same global features among all points, we project the extracted image features onto the partially denoised point cloud at each diffusion step. Therefore, we design a corresponding projection function $\pi:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{N\times C}$, where $N$ denotes the number of points in the point cloud. We implement $\pi$ using the efficient point cloud rasterizer [56]. For every individual point within the point cloud, we obtain semantic features corresponding to the pixel onto which the point is rasterized. Subsequently, the pointwise semantic features at timestep $t$ can be represented as $X_{t}^{O}=\pi(\mathcal{F},\phi_{C},X_{t})\in\mathbb{R}^{N\times C}$.
However, due to the inevitable occlusion induced by the hand, it is very challenging to reconstruct the occluded part of the object as only a single view is provided. Due to the incomplete object semantic features, relying solely on the local features of object points alone is insufficient to conduct precise hand-held object reconstruction. On the other side, the hand information should not be fully ignored as the hand naturally contains information on the shape of the hand-occluded object region. Therefore, we apply $\pi$ to both object and hand points to obtain $X_{t}^{HO}=\pi(\mathcal{F},\phi_{C},[X_{t},P_{H}])\in\mathbb{R}^{(N+N_{H})\times C}$, which serves as a unified semantic hand-object embedding. Thereby, $N_{H}$ denotes the number of the hand mesh vertices $P_{H}$, which can be obtained from the estimated hand pose $\phi_{H}$ together with the MANO hand model [59]. Noteworthy, we also apply an extra one-hot encoding to indicate whether the points belong to the object or the hand, hence making $X_{t}^{HO}\in\mathbb{R}^{(N+N_{H})\times(C+1)}$.
Geometric Hand Articulation Embedding. Furthermore, the object shape is also highly constrained by the hand and very geometrically related to the hand articulation. In the core, the positions and directions of the hand joints in contact with the object, are supposed to be in accordance with the object surface for a stable hold. Therefore, knowing these properties can serve as a strong prior to enhance reconstruction quality. In practice, we explicitly encode hand-object interaction by transforming the partially denoised point cloud at every step to each hand joint frame.
Given the partially denoised point cloud $X_{t}$ and hand parameter $\phi_{H}$, we first compute the rotation $R_{j}$ and translation $T_{j}$ of each joint $j$ with respective to the hand wrist using forward kinematics given the hand model. We then transform $X_{t}$ from the hand wrist coordinate system to each hand joint frame via $X_{t}^{j}=R_{j}X_{t}+T_{j}\in\mathbb{R}^{N\times 3}$ to encode the hand articulation onto object points. Finally, the hand articulation embedding $X_{t}^{A}\in\mathbb{R}^{N\times J}$ is calculated via concatenation and flattening of $X_{t}^{j}$, with $J$ equaling 45 in our experiments as 15 hand joints are utilized.
Dual-Stream Conditioning. One possible approach is to directly concatenate the unified semantic hand-object embedding $X_{t}^{HO}$ and geometric hand articulation embedding $X_{t}^{A}$ as $z=[X_{t}^{HO},X_{t}^{A}]$ and use it as condition. However, forcibly integrating semantic and geometric embeddings may lead to mutual interference. Instead, we propose a dual-stream conditioning strategy to fuse $X_{t}^{HO}$ and $X_{t}^{A}$. In essence, we process $z_{1}=X_{t}^{HO}$ and $z_{2}=X_{t}^{A}$ separately through two distinct network branches, resulting in two features of identical dimensions, as shown in Fig. 2. The noise $\epsilon_{\theta}$ is then predicted from the concatenated feature.
3.3 Hand-Constrained Centroid Fixing

### 18 Mar 2024 v2

# D-SCo: Dual-Stream Conditional Diffusion for Monocular Hand-Held Object Reconstruction

@@ Abstract
Reconstructing hand-held objects from a single RGB image is a challenging task in computer vision. In contrast to prior works that utilize deterministic modeling paradigms, we employ a point cloud denoising diffusion model to account for the probabilistic nature of this problem. In the core, we introduce centroid-fixed Dual-Stream Conditional diffusion for monocular hand-held object reconstruction (D-SCo), tackling two predominant challenges. First, to avoid the object centroid from deviating, we utilize a novel hand-constrained centroid fixing paradigm, enhancing the stability of diffusion and reverse processes and the precision of feature projection. Second, we introduce a dual-stream denoiser to semantically and geometrically model hand-object interactions with a novel unified hand-object semantic embedding, enhancing the reconstruction performance of the hand-occluded region of the object. Experiments on the synthetic ObMan dataset and three real-world datasets HO3D, MOW and DexYCB demonstrate that our approach can surpass all other state-of-the-art methods. Codes will be released.
Keywords:
Point Cloud Denoising Diffusion Model Hand-Held Object Reconstruction
@@ 1 Introduction
Figure 1: Comparison between D-SCo and naive diffusion models for hand-held object reconstruction. Naive diffusion models are conditioned only on image features without controlling object centroid deviation or modeling the uncertainty induced by hand occlusion. D-SCo, however, keeps the object centroid fixed under the constraint of the hand, making the diffusion model focus on shape reconstruction, and utilizes a dual-stream architecture to individually process semantic and geometric priors to learn a suitable representation for their own domain, tackling the aforementioned problems.
@@
Modeling hand-object interaction has a wide range of applications across various domains, including AR/VR [38, 55, 56] and human-robot interaction [20, 50, 23, 77]. Significant progress has been recently made for this task [38, 77, 4], where monocular hand-held object reconstruction draws particular attention [31, 37, 78]. Reconstructing hand-held objects from a single RGB image is a highly challenging and ill-posed problem. Suffering from the lack of real-world data and the ambiguity caused by hand- and self-occlusion, the performance of single-view hand-held object reconstruction remains limited [31, 78, 80].
@@
In essence, most existing works either rely on Signed Distance Fields (SDFs) [37, 78] or Directed Distance Fields (DDFs) [80] to represent object shapes. Despite being effective under favorable conditions, such techniques tend to result in over-smoothed and undetailed reconstruction [52, 11]. Moreover, prior work [31, 37, 78, 80] typically utilize a deterministic modeling paradigm, making it difficult to reason about the uncertainties introduced by hand- or self-occlusion. Recently, probabilistic point cloud denoising diffusion models [42, 83, 44] have shown to be effective for the task of single-view object reconstruction. Compared to works employing surface-based representations like SDFs and DDFs, diffusion-driven methods for reconstructing point clouds enjoy better capabilities in overcoming artifacts such as noise, sparsity, and irregularities [11]. Moreover, the probabilistic nature of denoising diffusion models is particularly beneficial for modeling uncertainties and underconstrained problems.
@@
Thus, in this work, we propose to leverage these probabilistic point cloud denoising diffusion models to conduct hand-held object reconstruction from a single RGB image. However, directly employing diffusion models in single-view hand-held object reconstruction faces two main problems: First, in existing diffusion models [43, 74, 75], the centroid of a partially denoised point cloud is not controlled and can thus deviate, for example, to the back side of the hand or even intersect with the hand, leading to physically implausible results. Additionally, the centroid deviation causes the misalignment of the semantic features [16] and can thus have adverse effects on the object reconstruction quality. Second, diffusion models for 3D reconstruction are typically conditioned only on single-stream 2D image features [74, 44], without modeling geometric hand-object interaction or addressing the uncertainty induced by hand occlusion.
@@
To solve the aforementioned problems, we present D-SCo, a centroid-fixed dual-stream conditional point cloud denoising diffusion model for single-view hand-held object reconstruction. As shown in Fig. 1, we compare our D-SCo with naive diffusion models. First, a hand-constrained centroid fixing scheme is proposed to ensure that the centroid of the partially denoised point cloud does not diverge during the diffusion as well as the reverse process. In particular, we use a small and efficient neural network to estimate the hand-constrained object centroid and use it as a guide during the reverse process. Hence, the model only needs to consider the simpler shape diffusion task rather than having to simultaneously account for shape and position. Second, to best leverage the hand-object interaction prior, we introduce a dual-stream denoiser, which individually processes semantic and geometric priors. In particular, we utilize a unified hand-object semantic embedding to compensate for hand occlusion in the 2D feature extraction stage.
Our contributions can be summarized as follows.
  • We present D-SCo, the first conditional point cloud diffusion model for 3D reconstruction of hand-held objects from a single RGB image.
@@
  • We introduce a novel hand-constrained centroid fixing scheme, utilizing the hand vertices prior to prevent the centroid of the partially denoised point cloud from diverging during diffusion and reverse processes.
  • We further propose a dual-stream denoiser to semantically and geometrically model hand-object interactions. Our novel unified hand-object semantic embedding serves as a strong prior for reconstructing occluded regions of the object.
@@
Experiments on the synthetic ObMan [31] dataset and three real-world datasets, i.e., HO3D [28], MOW [4] and DexYCB [6], demonstrate that D-SCo can surpass all existing methods by a large margin.
@@ 2 Related Work
Single-view Object Reconstruction. Object reconstruction is a long-standing problem in the computer vision community. Traditionally, varieties of works focus on reconstruction from multi-view geometric techniques [29, 64, 62, 63]. Recently, advanced by the powerful representation capacity of deep learning, single-view-based methods for object reconstruction have shown promising results [12, 74, 75], despite the highly ill-posed setting. Early learning-based methods propose to learn category-specific networks [36, 35, 70], while more recent works try to learn a generalizable model across multiple categories for either meshes [27, 25], voxels [24, 12, 73], point clouds [21, 39], or implicit representations such as NeRFs [45] and SDFs [52, 10]. In this work, we focus on the difficult yet very important problem of learning hand-held object reconstruction from a single RGB image [31], putting a particular emphasis on modeling the impact of the hand occlusion.
Single-View Hand-Held Object Reconstruction. Reconstructing hand-held objects is very challenging due to intricate hand- and self-occlusions. Earlier works opt for the simplified task of 6DoF object pose estimation [72, 67, 69, 28, 30], assuming known object templates. To better utilize hand-object interactions, some works jointly reason about hand and object poses by means of implicit feature fusion [7, 40, 65, 69], utilizing explicit geometric constraints [13, 2, 26, 4], or enforcing of physical constraints [71, 53]. Increasing attention has recently been paid to model-free hand-held object reconstruction, as it is a more applicable setting. Exemplarily, while [31] explicitly predicts the object mesh, [37, 9, 8, 80] leverage implicit 3D representations such as SDFs, DDFs for reconstructing the shape of the hand and object. Recently, the concurrent work [81] instead utilizes NeRF as a representation. However, all these methods follow a deterministic approach, oftentimes resulting in low-quality reconstruction results for occluded and invisible parts, especially in the ill-posed single-view setting. We instead draw inspiration from recent advances in probabilistic 3D generation [33, 83, 16, 44] and additionally leverage semantic and geometric hand-object priors to achieve object reconstructions with high fidelity.
@@
Hand Pose estimation. Hand pose estimation methods from RGB(-D) images can be primarily categorized into model-free and model-based methods. Model-free approaches typically detect 2D keypoints, which are then lifted to 3D joint positions [58, 59, 86, 47, 51, 34, 48], whereas model-based approaches commonly exploit statistical models such as MANO [60] to work in a low-dimensional parameter spaces [66, 1, 82, 85, 61]. Compared with model-free methods, the latter line of works exhibits better robustness to occlusions as well as domain discrepancies [78]. Consequently, in this work, we rely on an off-the-shelf model-based approach [31, 61] to obtain the hand poses, which are then further leveraged within object shape inference.
@@ 3 Method
Figure 2: Architecture of D-SCo. (I) Given a single-view RGB image, we first predict the hand pose $\phi_{H}$ and camera view $\phi_{C}$ by an off-the-shelf network. (II) The object centroid $\widehat{\mathcal{M}}$ is then estimated by our simple yet efficient hand-constrained centroid prediction network. (III) We further introduce a centroid-fixed diffusion network, which always keeps the centroid of partially denoised point cloud fixed at the predicted centroid $\widehat{\mathcal{M}}$ during the reverse process. (IV) A dual-stream denoiser is proposed to individually process and then aggregate semantic and geometric hand-object interaction priors as condition. A unified hand-object semantic embedding is introduced to serve as a strong prior of hand-occlusion.
Given a single RGB image $I$ capturing a hand-held object, we aim at reconstructing its shape as a 3D point cloud. As shown in Fig. 2, we first utilize off-the-shelf methods [31, 61] to predict hand parameters $\phi_{H}$ and camera view $\phi_{C}$. Thereby, $\phi_{H}$ is defined with respect to the MANO [60] hand model, having 45DoF joint parameters, and $\phi_{C}$ represents the 6DoF pose of the hand wrist in the world reference system. Given the estimated hand vertices as obtained from the hand parameters, we employ a small yet efficient network to predict the object centroid $\widehat{\mathcal{M}}$. Eventually, we leverage the estimated centroid as part of the point cloud diffusion process to enable robust object point cloud reconstruction. Note that in contrast to existing diffusion methods, our model can thus fully focus on shape reconstruction, leading to improved performance. Furthermore, we introduce a dual-stream denoiser to first independently process and then aggregate semantic and geometric priors with a novel unified hand-object embedding, which helps the reconstruction of the hand-occluded regions of the object. Due to the probabilistic nature of diffusion models together with our explicit modeling of hand-object interaction, D-SCo shows superior performance in handling uncertainties arising from hand- and self-occlusion.
In the following sections, we first introduce the fundamentals of conditional point cloud denoising diffusion models (Sec. 3.1). We then explain our centroid-fixed conditional diffusion (Sec. 3.2) before diving into our proposed dual-stream conditional denoiser (Sec. 3.3).
@@ 3.1 Conditional Point Cloud Denoising Diffusion
We formulate single-view hand-held object reconstruction as conditional point cloud denoising diffusion, consisting of two Markov chains called the diffusion process and the reverse diffusion process. Suppose that we have a target point cloud with $N$ points $X_{0}\in\mathbb{R}^{3N}$ from the conditional distribution $q(X|z)$, where $z=I_{\phi_{C},\phi_{H}}$ is the input RGB image with the corresponding camera view $\phi_{C}$ and the hand pose $\phi_{H}$. For the diffusion process, we gradually add Gaussian noise to the target point cloud at different levels $t\in\{1,\cdots,T\}$ as
@@
Starting with a point cloud sample $X_{T}$ from random Gaussian noise, the reverse process iteratively samples from $q(X_{t-1}|X_{t},z)$ to remove the added noise from the diffusion process. To approximate this reverse process, we train a point cloud conditional denoiser $\mathcal{D}_{\theta}(X_{t},t,z)$ to learn the distribution $q(X_{t-1}|X_{t},z)$ by
@@
3.2 Hand-Constrained Centroid-Fixed Conditional Diffusion